# Anisotropy of quadratic forms over global fields of characteristic $\neq$ 2 is Diophantine

### v2

@@ 4.1. Dealing with a Single Prime
  • (2)
    $\{(x,y)\in K^{\times}\times K^{\times}:(x,y)_{\mathfrak{p}}=-1\}$.
If $|\mathbb{F}_{\mathfrak{p}}|$ is odd, then it follows from Lemma 4.1 and Lemma 4.3 that (2) and (3) are diophantine sets since $\mathcal{O}_{\mathfrak{p}}$ and $\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$ are diophantine over $K$.
If $K$ is a number field and $\mathfrak{p}|2$ then it was stated in the proof of Theorem 4.5 of [EM18] that $K_{\mathfrak{p}}^{\times}:{K_{\mathfrak{p}}^{\times}}^{2}]$ is finite (see also 63:9 in [O’M00]). Let $s_{1},...s_{n}\in K$ be a set of representatives for ${K_{\mathfrak{p}}}^{\times}/{{K_{\mathfrak{p}}}^{\times}}^{2}$ and $S_{i}:=s_{j}\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})$ where $e$ is the absolute ramification index. Then $x\in{K_{\mathfrak{p}}^{\times}}^{2}$ if and only if
$x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}\vee x\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}}).$
$(x,y)_{\mathfrak{p}}=1$ if and only if
$(x,y)\in\bigcup_{i,j:(s_{i},s_{j})_{\mathfrak{p}}=1}S_{i}\times S_{j}.$
$(x,y)_{\mathfrak{p}}=-1$ if and only if
$(x,y)\in\bigcup_{i,j:(s_{i},s_{j})_{\mathfrak{p}}=-1}S_{i}\times S_{j}.$
For a real infinite prime $\mathfrak{p}$, Let $\omega:K\to\mathbb{R}$ be the corresponding embedding of $K$ in to $K_{\mathfrak{p}}=\mathbb{R}$. The Lemma 4.1 of [EM18] say that the set $\{x\in K^{\times}:\omega(x)>0\}$ is diophantine over $K$. Then $f$ is anisotropic if and only if
$(\omega(a_{1})>0\wedge\omega(a_{2})>0\wedge\omega(a_{3})>0\wedge\omega(a_{4})>0)\vee(\omega(a_{1})<0\wedge\omega(a_{2})<0\wedge\omega(a_{3})<0\wedge\omega(a_{4})<0).$
Therefore, these sets are all diophantine sets.
@@ 4.2. The proof of main theorem
Now we can prove our main theorem: Theorem 1.2.
@@
For $\sigma\in\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)$ with $\sigma\neq(1,1)$, let $s_{\sigma}=a$ if $\sigma=(-1,\pm 1)$ and $s_{\sigma}=b$ if $\sigma=(1,-1)$.
We claim that $f\textup{ is anisotropic over some }\mathfrak{p}\in S_{K}$ if and only if one of the following conditions holds:
  • $\exists\mathfrak{p}|\mathfrak{m}$ such that $f$ is anisotropic over $K_{\mathfrak{p}}$.
  • $\bigvee_{\sigma\neq(1,1)}\quad\exists p\in\Phi_{\sigma}$ such that $(\dagger_{a_{1},a_{2}}^{+}\wedge\dagger_{-a_{3},-a_{4}}^{-})\vee(\dagger_{a_{1},a_{2}}^{-}\wedge\dagger_{-a_{3},-a_{4}}^{+})\wedge\mathfrak{e}$
    where the formula $\mathfrak{e}$ is $a_{1}a_{2}a_{3}a_{4}\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times}\wedge a_{1}a_{2}a_{3}a_{4}\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))$.
    and a formula $\dagger_{x,y}^{-}$ is defined by
    $((x\in p\cdot{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\wedge(y\textup{ or }-xy\in s_{\sigma}\cdot{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
    $\vee((y\in p\cdot{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\wedge(x\textup{ or }-yx\in s_{\sigma}\cdot{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
    and a formula $\dagger_{x,y}^{+}$ is defined by
    $((x\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
    $\wedge((y\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(x\textup{ and }-yx\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
  • $\exists(p,q)\in\Psi_{K}$ such that $q\in(R_{p,q}^{(1,1)})^{\times}$ and $(\ddagger_{a_{1},a_{2}}^{+}\wedge\ddagger_{-a_{3},-a_{4}}^{-})\vee(\ddagger_{a_{1},a_{2}}^{-}\wedge\ddagger_{-a_{3},-a_{4}}^{+})\wedge\mathfrak{e}$,
    where a formula $\ddagger_{x,y}^{-}$ is defined by
    $((x\in p\cdot{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\wedge(y\textup{ or }-xy\in q\cdot{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
    $\vee((y\in p\cdot{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\wedge(x\textup{ or }-yx\in q\cdot{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
    and a formula $\ddagger_{x,y}^{+}$ is defined by
    $((x\in{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
    $\wedge((y\in{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\vee(x\textup{ and }-yx\in{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
For the first situation, it gives a diophantine set since there are at most finitely many primes in $\mathfrak{m}$, as well as Theorem 4.4.
The second and third situation also give diophantine sets by Proposition 3.9(1) as well as Remark 3.10.
As a conclusion, the set
is a diophantine set, as a union of that three conditions.
It is left for us to show that the claim holds.
@@
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\textup{ and }a_{1}a_{2}a_{3}a_{4}\in(K_{p}^{\times})^{2}.$
So it is sufficient to prove the following two equivalent relations between the statements in our claim and this condition for different types of $\mathfrak{p}$.
@@
    $\exists p\in\Phi_{\sigma}\textup{ such that }\dagger\wedge\mathfrak{e}$
    is equivalent to the condition that there exists a prime $\mathfrak{p}$ with $\psi_{L/K}(\mathfrak{p})=\sigma$ such that $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}$.
@@
    $\exists(p,q)\in\Psi_{K}\textup{ such that }q\in(R_{p,q}^{\sigma})^{\times}\textup{ and }\ddagger\wedge\mathfrak{e}$
    is equivalent to the condition that there exists a prime $\mathfrak{p}$ with $\psi_{L/K}(\mathfrak{p})=(1,1)$ such that $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in(K_{p}^{\times})^{2}$.
@@
For $\sigma\neq(1,1)$, if there exists $p\in\Phi_{\sigma}$ such that $\dagger$ holds, then by Proposition 3.9(2) we know that $\mathbb{P}^{\sigma}(p)$ is nonempty. For $\mathfrak{p}\in\mathbb{P}^{\sigma}(p)$, recall that the formula $\dagger_{x,y}^{-}$ is
$((x\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
$\wedge((y\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(x\textup{ and }-yx\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
@@
By Lemma 3.4, and the definition of $R_{\mathfrak{p}}^{\sigma}$, we have $R_{p}^{\sigma}\subset\mathcal{O}_{\mathfrak{p}}$. Also by Proposition 3.6 we have $J(R_{p}^{\sigma})\subset\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$. Hence $\dagger_{x,y}^{-}$ implies $(x,y)_{\mathfrak{p}}=-1$ by $v_{\mathfrak{p}}(p)=1$, Lemma 4.1 and the fact that $s_{\sigma}$ is a $\mathfrak{p}$-adic unit which is not a square by the construction of $s_{\sigma}$ and the definition of Artin map. For the same reason, $\dagger_{x,y}^{+}$ implies $(x,y)_{\mathfrak{p}}=1$, by Lemma 4.2. The formula $\mathfrak{e}$ implies
$a_{1}a_{2}a_{3}a_{3}\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}\cap{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}),$
which means $a_{1}a_{2}a_{3}a_{4}$ is a square in $\mathcal{O}_{\mathfrak{p}}$. So $\dagger\wedge\mathfrak{e}$ will imply $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$.
@@
Conversely, if there exists a prime $\mathfrak{p}\nmid\mathfrak{m}$ with $\psi_{L/K}({\mathfrak{p}})=\sigma$ such that $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$, then by Proposition 3.9(3) we can find $p\in\Phi_{\sigma}$ such that $\mathbb{P}^{\sigma}(p)=\{\mathfrak{p}\}$. In this situation we have $R_{p}^{\sigma}=\mathcal{O}_{\mathfrak{p}}$ for $\sigma\neq(1,1)$. So the formula $\dagger_{x,y}^{\pm}$ can be deduced from $(x,y)_{\mathfrak{p}}=\pm 1$. Hence, the formula $\dagger\wedge\mathfrak{e}$ can be deduced from $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$.
A similar reason can be used to obtain the second equivalent relation, by using Proposition 3.9(4) and (5). Now we specify the process of proof.
Assume $\mathfrak{p}_{0}\in I_{\mathfrak{m}}$ with $\Psi_{L/K}(\mathfrak{p})=(1,1)$. By Lemma 3.11 and Proposition 3.9, we can find $(p,q)\in\Psi_{K}$ with $\Delta_{ap,q}\cap\Delta_{bp,q}=\{\mathfrak{p}_{0}\}$. Moreover, we have $v_{\mathfrak{p}_{0}}(p)=1$, $v_{\mathfrak{p}_{0}}(q)=0$ and $q$ is not a square modulo $\mathfrak{p}$ by Lemma 3.11. Then by Lemma 4.1, Lemma 4.2 and Lemma 4.3, $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in{K_{\mathfrak{p}}^{\times}}^{2}$ implies $\ddagger\wedge\mathfrak{e}$ since $R_{p,q}^{(1,1)}=\mathcal{O}_{\mathfrak{p}}$ and $J(R_{p,q}^{(1,1)})=\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$.
@@
Conversely, if there exists $(p,q)\in\Psi_{K}$, then by Proposition 3.9, $\Delta_{ap,q}\cap\Delta_{bp,q}\neq\emptyset$. For $\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}$ since $q\in({R_{p,q}^{1,1}})^{\times}$ and $(ap,q)_{\mathfrak{p}}=(bp,q)_{\mathfrak{p}}=-1$, we have
$\big(\frac{q}{\mathfrak{p}}\big)^{v_{\mathfrak{p}}(bp)}=\big(\frac{q}{\mathfrak{p}}\big)^{v_{\mathfrak{p}}(ap)}=-1.$
So $q$ is not a square mod $\mathfrak{p}$ and $v_{\mathfrak{p}}(ap),v_{\mathfrak{p}}(bp)$ are odd. Since the support of $a$ and $b$ are disjoint (Lemma 3.4), we get $v_{\mathfrak{p}}(p)$ is odd. Again by Lemma 4.1, Lemma 4.2 and Lemma 4.3, $\ddagger\wedge\mathfrak{e}$ implies $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in{K_{\mathfrak{p}}^{\times}}^{2}$ since $R_{p,q}^{(1,1)}\subset\mathcal{O}_{\mathfrak{p}}$ and $J(R_{p,q}^{(1,1)})\subset\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$. ∎
@@
    $\{(a_{1},a_{2})\in{K^{\times}}\times K^{\times}:-(a_{1}a_{2})\notin{K^{\times}}^{2}\}$
@@
    $a_{3}=a_{1}(x_{1}^{2}+a_{1}a_{2}(\frac{x_{2}}{a_{1}})^{2})\in a_{1}\textup{N}_{K(\sqrt{a_{1}a_{2}})/K}(K(\sqrt{a_{1}a_{2}})^{\times}).$
    Hence for all $a_{1},a_{2},a_{3}\neq 0$, $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{3}$ is isotropic if and only if $a_{1}a_{2}\in{K^{\times}}^{2}$ or $a_{3}\in a_{1}\cdot\textup{N}_{K(\sqrt{a_{1}a_{2}})/K}(K(\sqrt{a_{1}a_{2}})^{\times})$.
@@
    $\{(a_{1},a_{2},a_{3})\in{K^{\times}}\times K^{\times}\times K^{\times}:a_{1}a_{2}\neq K^{2})\}$

### v3

@@ 4.1. Dealing with a Single Prime
  3. (3)
    $\{(x,y)\in K^{\times}\times K^{\times}:(x,y)_{\mathfrak{p}}=-1\}$
Based on the first three chapters of [RUM80], the valuation ring $\mathcal{O}_{\mathfrak{p}}$ is diophantine over $K$. Consequently, the maximal ideal $\mathfrak{p}\mathcal{O}_{\mathfrak{p}}=\pi\mathcal{O}_{\mathfrak{p}}$ (where $v_{\mathfrak{p}}(\pi)=1$) is also diophantine. More precisely, Theorem 1 of [RUM80] implies that $\mathcal{O}_{\mathfrak{p}}$ is definable, and since the predicate $S_{l}(x,\vec{c})$ is given by an existential positive formula, it is diophantine.
If $|\mathbb{F}_{\mathfrak{p}}|$ is odd, the diophantine nature of (2) and (3) follows from Lemma 4.1 and Lemma 4.3, as these sets are constructed from $\mathcal{O}_{\mathfrak{p}}$ and $\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$ using field operations and logical disjunction/conjunction. Additionally, Lemma 4.2 (3) ensures that the set in (1) is diophantine.
If $K$ is a number field and $\mathfrak{p}|2$, the index $K_{\mathfrak{p}}^{\times}:{K_{\mathfrak{p}}^{\times}}^{2}]$ is finite (see [O’M00], 63:9). Let $s_{1},\ldots,s_{n}\in K^{\times}$ be a set of representatives for the cosets of ${K_{\mathfrak{p}}^{\times}}^{2}$ in $K_{\mathfrak{p}}^{\times}$. Such representatives can be chosen in $K^{\times}$ because $K$ is dense in $K_{\mathfrak{p}}$ and the cosets are open in the local topology. Let $e$ be the absolute ramification index. We define the sets
$S_{i}:=s_{i}\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})\quad\text{for }i=1,\ldots,n.$
These sets $S_{1},\ldots,S_{n}$ form a partition of $K^{\times}\cap K_{\mathfrak{p}}^{\times}$. This follows from two facts:
  • The sets are disjoint: if $i\neq j$, then $s_{i}s_{j}^{-1}\notin{K_{\mathfrak{p}}^{\times}}^{2}$. By Lemma 4.4, the set ${K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})$, being a subset of ${K_{\mathfrak{p}}^{\times}}^{2}$, cannot contain $s_{i}s_{j}^{-1}$.
  • They cover the relevant elements: suppose $x\in K\cap s_{i}{K_{\mathfrak{p}}^{\times}}^{2}$, then $s_{i}^{-1}x\in K\cap{K_{\mathfrak{p}}^{\times}}^{2}$. By Lemma 4.4, we have $s_{i}^{-1}x\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})$, which implies $x\in S_{i}$.
Under this partition, $x\in{K_{\mathfrak{p}}^{\times}}^{2}$ if and only if $x\in S_{i}$ for $s_{i}\in{K_{\mathfrak{p}}^{\times}}^{2}$. Furthermore, the Hilbert symbol conditions satisfy:
$(x,y)_{\mathfrak{p}}=1\iff(x,y)\in\bigcup_{(s_{i},s_{j})_{\mathfrak{p}}=1}(S_{i}\times S_{j}),$
$(x,y)_{\mathfrak{p}}=-1\iff(x,y)\in\bigcup_{(s_{i},s_{j})_{\mathfrak{p}}=-1}(S_{i}\times S_{j}).$
Since $\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}}$ is diophantine, each $S_{i}$ is diophantine, and so are their finite unions.
For a real infinite prime $\mathfrak{p}$, let $\omega\colon K\hookrightarrow\mathbb{R}$ be the corresponding embedding. Lemma 4.1 of [EM18] states that the set $P_{\omega}:=\{x\in K^{\times}:\omega(x)>0\}$ is diophantine over $K$. The form $f$ is anisotropic over $K_{\mathfrak{p}}\cong\mathbb{R}$ if and only if it is definite, i.e.,
$\left(\bigwedge_{i=1}^{4}a_{i}\in P_{\omega}\right)\vee\left(\bigwedge_{i=1}^{4}-a_{i}\in P_{\omega}\right).$
Since $P_{\omega}$ is diophantine, this condition defines a diophantine set. ∎
@@ 4.2. The proof of main theorem
Now we prove our main result, Theorem 1.3.
@@
For $\sigma\in\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)$ with $\sigma\neq(1,1)$, define
$s_{\sigma}:=\begin{cases}a,&\text{if }\sigma=(-1,\pm 1),\\ b,&\text{if }\sigma=(1,-1).\end{cases}$
We claim that the quadratic form
$f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}$
is anisotropic over some $\mathfrak{p}\in S_{K}$ if and only if one of the following three conditions holds.
  • (i)
    There exists a prime $\mathfrak{p}\mid\mathfrak{m}$ such that $f$ is anisotropic over $K_{\mathfrak{p}}$.
  • (ii)
    There exists $\sigma\neq(1,1)$ and an element $p\in\Phi_{\sigma}$ such that
    $\Bigl(\dagger^{+}_{a_{1},a_{2}}\wedge\dagger^{-}_{-a_{3},-a_{4}}\Bigr)\;\vee\;\Bigl(\dagger^{-}_{a_{1},a_{2}}\wedge\dagger^{+}_{-a_{3},-a_{4}}\Bigr)\;\wedge\;\mathfrak{e},$
    where
    $\mathfrak{e}\;:\;a_{1}a_{2}a_{3}a_{4}\in K^{\times 2}\cdot(1+J(R_{p}^{\sigma})).$
    Here the formulas $\dagger^{\pm}_{x,y}$ are defined as follows.
@@
[…]
  • (iii)
    There exists $(p,q)\in\Psi_{K}$ such that $q\in(R_{p,q}^{(1,1)})^{\times}$ and
    $\Bigl(\ddagger^{+}_{a_{1},a_{2}}\wedge\ddagger^{-}_{-a_{3},-a_{4}}\Bigr)\;\vee\;\Bigl(\ddagger^{-}_{a_{1},a_{2}}\wedge\ddagger^{+}_{-a_{3},-a_{4}}\Bigr)\;\wedge\;\mathfrak{e}^{\prime},$
    where $\mathfrak{e}^{\prime}$ is defined analogously to $\mathfrak{e}$, with $R_{p}^{\sigma}$ replaced by $R_{p,q}^{(1,1)}$.
    The formulas $\ddagger^{\pm}_{x,y}$ are given by
[…]
The first condition defines a diophantine set, since the set $\mathfrak{m}$ contains only finitely many primes and by Theorem 4.5. The second and third conditions also define diophantine sets by Proposition 3.9(1) together with Remark 3.10. Consequently, the set
$\bigl\{(a_{1},a_{2},a_{3},a_{4})\in(K^{\times})^{4}\mid f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}\textup{ is anisotropic over some }\mathfrak{p}\in S_{K}\bigr\}$
is a diophantine set, being a finite union of the three sets defined above.
It remains to verify that the claim holds. For brevity, we write
[…]
and
[…]
The primes $\mathfrak{p}\nmid\mathfrak{m}$ can be classified according to $\sigma\in\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)$. By Theorem 2.3(3), the condition that $f$ is anisotropic over $\mathfrak{p}\nmid\mathfrak{m}$ is equivalent to
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
Hence, it suffices to establish the following two equivalences between the statements in our claim and this local condition for the different types of primes $\mathfrak{p}$.
@@
    $\exists p\in\Phi_{\sigma}\text{ such that }\dagger\wedge\mathfrak{e}$
    is equivalent to the existence of a prime $\mathfrak{p}$ with $\psi_{L/K}(\mathfrak{p})=\sigma$ such that
    $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
@@
    $\exists(p,q)\in\Psi_{K}\text{ such that }q\in(R_{p,q}^{(1,1)})^{\times}\text{ and }\ddagger\wedge\mathfrak{e}^{\prime}$
    is equivalent to the existence of a prime $\mathfrak{p}$ with $\psi_{L/K}(\mathfrak{p})=(1,1)$ such that
    $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
@@
For $\sigma\neq(1,1)$, suppose there exists $p\in\Phi_{\sigma}$ such that $\dagger$ holds. Then, by Proposition 3.9(2), the set $\mathbb{P}^{\sigma}(p)$ is nonempty.
Let $\mathfrak{p}\in\mathbb{P}^{\sigma}(p)$. Recall that the formula $\dagger_{x,y}^{-}$ is given by
$((x\in p\cdot K^{\times 2}\cdot(R_{p}^{\sigma})^{\times})\wedge(y\text{ or }-xy\in s_{\sigma}\cdot K^{\times 2}\cdot(1+J(R_{p}^{\sigma}))))$
$\vee((y\in p\cdot K^{\times 2}\cdot(R_{p}^{\sigma})^{\times})\wedge(x\text{ or }-yx\in s_{\sigma}\cdot K^{\times 2}\cdot(1+J(R_{p}^{\sigma})))).$
By Lemma 3.4 and the definition of $R_{p}^{\sigma}$, we have $R_{p}^{\sigma}\subset\mathcal{O}_{\mathfrak{p}}$. Moreover, by Proposition 3.6, $J(R_{p}^{\sigma})\subset\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$.
@@
Hence, $\dagger_{x,y}^{-}$ implies $(x,y)_{\mathfrak{p}}=-1$, because $v_{\mathfrak{p}}(p)$ is odd, by Lemma 4.1, and since $s_{\sigma}$ is a $\mathfrak{p}$-adic unit which is not a square (by the construction of $s_{\sigma}$ and the definition of the Artin map). For example, if $\sigma=(1,-1)$, then $b$ is not a square modulo $\mathfrak{p}$ since $\psi_{L/K}(\mathfrak{p})=(1,-1)$.
By the same reasoning, $\dagger_{x,y}^{+}$ implies $(x,y)_{\mathfrak{p}}=1$.
Finally, the formula $\mathfrak{e}$ implies
$a_{1}a_{2}a_{3}a_{4}\in K^{\times 2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}),$
which means that $a_{1}a_{2}a_{3}a_{4}$ is a square in $K_{\mathfrak{p}}$ by Lemma 4.2(3).
Therefore, $\dagger\wedge\mathfrak{e}$ implies
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
@@
Conversely, suppose there exists a prime $\mathfrak{p}\nmid\mathfrak{m}$ with $\psi_{L/K}(\mathfrak{p})=\sigma$ such that
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}.$
Then, by Proposition 3.9(3), we can find $p\in\Phi_{\sigma}$ such that $\mathbb{P}^{\sigma}(p)=\{\mathfrak{p}\}$. In this situation, for $\sigma\neq(1,1)$, we have
$R_{p}^{\sigma}=\mathcal{O}_{\mathfrak{p}}.$
Therefore, the formula $\dagger_{x,y}^{\pm}$ can be deduced directly from $(x,y)_{\mathfrak{p}}=\pm 1$, and the formula $\mathfrak{e}$ can be deduced from
$a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
Hence, we conclude that
$\dagger\wedge\mathfrak{e}$
follows from
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
Conversely, suppose there exists a prime $\mathfrak{p}\nmid\mathfrak{m}$ with $\psi_{L/K}(\mathfrak{p})=\sigma$ such that
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}.$
Then, by Proposition 3.9(3), we can find $p\in\Phi_{\sigma}$ such that $\mathbb{P}^{\sigma}(p)=\{\mathfrak{p}\}$. In this situation, for $\sigma\neq(1,1)$, we have
$R_{p}^{\sigma}=\mathcal{O}_{\mathfrak{p}}.$
Therefore, the formula $\dagger_{x,y}^{\pm}$ can be deduced directly from $(x,y)_{\mathfrak{p}}=\pm 1$, and the formula $\mathfrak{e}$ can be deduced from
$a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
Hence, we conclude that
$\dagger\wedge\mathfrak{e}$
follows from
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}.$
A similar argument establishes the second equivalence, using Proposition 3.9(4) and (5). We now describe the proof for the case $\psi_{L/K}(\mathfrak{p})=(1,1)$.
Assume $\mathfrak{p}_{0}\in I_{\mathfrak{m}}$ with $\Psi_{L/K}(\mathfrak{p}_{0})=(1,1)$, then by Proposition 3.9, we can find $(p,q)\in\Psi_{K}$ such that
$\Delta_{ap,q}\cap\Delta_{bp,q}=\{\mathfrak{p}_{0}\}.$
Moreover, we have
$v_{\mathfrak{p}_{0}}(p)=1,\quad v_{\mathfrak{p}_{0}}(q)=0,$
and $q$ is not a square modulo $\mathfrak{p}_{0}$. Then, by Lemma 4.1, Lemma 4.2, and Lemma 4.3, the condition
$(a_{1},a_{2})_{\mathfrak{p}_{0}}=-(-a_{3},-a_{4})_{\mathfrak{p}_{0}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}_{0}}^{\times})^{2}$
implies
$\ddagger\wedge\mathfrak{e},$
since $R_{p,q}^{(1,1)}=\mathcal{O}_{\mathfrak{p}_{0}}$ and $J(R_{p,q}^{(1,1)})=\mathfrak{p}_{0}\mathcal{O}_{\mathfrak{p}_{0}}$.
@@
Conversely, suppose $(p,q)\in\Psi_{K}$, then by Proposition 3.9, we have $\Delta_{ap,q}\cap\Delta_{bp,q}\neq\emptyset$. For any $\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}$, since $q\in(R_{p,q}^{(1,1)})^{\times}$ and
$(ap,q)_{\mathfrak{p}}=(bp,q)_{\mathfrak{p}}=-1,$
we obtain
$\left(\frac{q}{\mathfrak{p}}\right)^{v_{\mathfrak{p}}(bp)}=\left(\frac{q}{\mathfrak{p}}\right)^{v_{\mathfrak{p}}(ap)}=-1.$
Hence $q$ is not a square modulo $\mathfrak{p}$, and $v_{\mathfrak{p}}(ap),v_{\mathfrak{p}}(bp)$ are odd. Since the supports of $a$ and $b$ are disjoint (Lemma 3.4), it follows that $v_{\mathfrak{p}}(p)$ is odd.
Again, by Lemma 4.1, Lemma 4.2, and Lemma 4.3, we conclude that
$\ddagger\wedge\mathfrak{e}\implies(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}\quad\text{and}\quad a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2},$
since $R_{p,q}^{(1,1)}\subset\mathcal{O}_{\mathfrak{p}}$ and $J(R_{p,q}^{(1,1)})\subset\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$. ∎
@@
    $\{(a_{1},a_{2})\in{K^{\times}}\times K^{\times}:-a_{1}a_{2}\notin{K^{\times}}^{2}\}$
@@
    $a_{3}=a_{1}(x_{1}^{2}+a_{1}a_{2}(\frac{x_{2}}{a_{1}})^{2})\in a_{1}\textup{N}_{K(\sqrt{-a_{1}a_{2}})/K}(K(\sqrt{-a_{1}a_{2}})^{\times}).$
    Hence for all $a_{1},a_{2},a_{3}\neq 0$, $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{3}$ is isotropic if and only if $a_{1}a_{2}\in{K^{\times}}^{2}$ or $a_{3}\in a_{1}\cdot\textup{N}_{K(\sqrt{-a_{1}a_{2}})/K}(K(\sqrt{-a_{1}a_{2}})^{\times})$.
@@
    $\{(a_{1},a_{2},a_{3})\in{K^{\times}}\times K^{\times}\times K^{\times}:-a_{1}a_{2}\notin{K^{\times}}^{2})\}$